# Slot: Provenance-Driven APT Detection through Graph Reinforcement Learning

### v2

@@ Abstract.
Advanced Persistent Threats (APTs) represent sophisticated cyberattacks characterized by their ability to remain undetected within the victim system for extended periods, aiming to exfiltrate sensitive data or disrupt operations. Existing detection approaches often struggle to effectively identify these complex threats, construct the attack chain for defense facilitation, or resist adversarial attacks. To overcome these challenges, we propose Slot, an advanced APT detection approach based on provenance graphs and graph reinforcement learning. Slot excels in uncovering multi-level hidden relationships, such as causal, contextual, and indirect connections, among system behaviors through provenance graph mining. By pioneering the integration of graph reinforcement learning, Slot dynamically adapts to new user activities and evolving attack strategies, enhancing its resilience against adversarial attacks. Additionally, Slot automatically constructs the attack chain according to detected attacks with clustering algorithms, providing precise identification of attack paths and facilitating the development of defense strategies. Evaluations with real-world datasets demonstrate Slot’s outstanding accuracy, efficiency, adaptability, and robustness in APT detection, with most metrics surpassing state-of-the-art methods. Additionally, case studies conducted to assess Slot’s effectiveness in supporting APT defense further establish it as a practical and reliable tool for cybersecurity protection.
@@ 1. Introduction
To effectively combat APTs, it is crucial to detect them early with both timeliness and precision. Traditional detection methods may struggle with these complex and prolonged attacks, often resulting in either inadequate accuracy or excessive time to produce detection results. Recently, researchers have increasingly utilized provenance graphs to trace APT activities, progressively achieving practical efficacy and efficiency.
@@
To address the aforementioned gaps, we introduce Slot, an advanced APT detection approach based on provenance graphs and graph reinforcement learning. Unlike existing methods, Slot offers improved accuracy in detecting APTs by thoroughly digging multi-level hidden relationships—such as causal, contextual, and indirect connections—within system behaviors using provenance graphs. Additionally, by incorporating graph reinforcement learning, Slot can autonomously learn and adapt to new user activities and attack strategies without relying on threat reports derived from expert insights, thereby enhancing its adaptability and resilience against adversarial attacks. Furthermore, Slot can automatically construct the attack chain, accurately identifying the attack path, which simplifies the defense process.
@@
We introduce Slot, an accurate APT detection method designed to withstand adversarial environments. It excels at identifying malicious activities within extensive audit logs and can also reconstruct the attack chain, thereby aiding in the formulation of effective defense strategies.
@@ 2.2. Graph Reinforcement Learning
Reinforcement Learning (RL) has recently achieved success in addressing challenges across various fields, including robotics (Levine et al., 2018), gaming (Chen et al., 2022), and natural language processing (NLP) (Uc-Cetina et al., 2023). Researchers have found that RL methods enable effective exploration of the topological structures and attribute information of graphs by analyzing key components such as nodes, links, and subgraphs. SUGAR (Sun et al., 2021) employs Q-learning to adaptively select significant subgraphs to represent discriminative information of the graphs. Bachu et al. (Bacciu and Numeroso, 2022) develop perturbation strategies for local explanations of graph data by optimizing multi-objective scores. Other related works (Yuan et al., 2021; Jiang et al., 2018) have demonstrated impressive performances in graph representation learning through RL, proving its efficacy in this domain. But currently, graph reinforcement learning has not yet been attempted in the field of log provenance detection.
@@ 4.2. Graph Construction
Slot constructs a comprehensive system provenance graph using audit data collected from logging infrastructures such as Windows ETW (Corporation, [n. d.]), Linux Audit (lin, [n. d.]), and CamFlow (Pasquier et al., 2018). It adheres to the existing definitions of provenance graphs and employs recent graph preprocessing techniques (Wang et al., 2022; Rehman et al., 2024), abstracting events into subjects, objects, and relationships, as illustrated in Table 2. Here, source nodes like Process/Thread represent subjects, and target nodes like Files or sockets represent objects, with system call events serving as the relationships to form graph . In addition to focusing on system behaviors, Slot further enriches the relationships between behaviors and explores multi-hop latent relations in Section 4.3. Regarding node features, Slot fully considers meaningful semantic information such as process names and command-line parameters for process nodes, file paths for file nodes, network IP addresses and ports for netflow nodes. By employing the Word2Vec model (Mikolov, 2013), Slot learns vector representations of these semantics, effectively mining the semantic expressions in provenance logs.
@@ 4.4.2. Adaptive Bandit Neighbor Selector
Reward The optimal at layer under relation aims to find the most similar neighbors to the central node. We design a binary reward based on the difference in average distances between two consecutive epochs. We define the reward for epoch as:
@@ 5.1.1. Datasets
The evaluation of Slot was carried out using four open-source datasets: StreamSpot (The, [n. d.]), Unicorn Wget (Han et al., 2020) and DARPA E3 (DAR, [n. d.]). Both StreamSpot and Unicorn consist of batch audit logs generated in controlled environments. The StreamSpot dataset includes 600 batch logs monitoring system calls across six distinct scenarios, five of which simulate benign user behavior, while one represents a drive-by download attack. The Unicorn dataset contains 150 batch logs collected via Camflow (Pasquier et al., 2018), with 125 logs corresponding to benign activity and 25 involving supply chain attacks. These attacks are stealthy in nature, designed to mimic normal system workflows, and are challenging to detect. The DARPA dataset is part of the DARPA Transparent Computing program, gathered during adversarial engagements in enterprise networks. The Red Team conducted APT attacks using various vulnerabilities to exfiltrate sensitive information, while the Blue Team focused on identifying these attacks through network host audits and causal analysis. We conducted experimental comparisons using datasets commonly shared across multiple SOTA algorithms (Wang et al., 2022; Rehman et al., 2024; Jia et al., 2024; Zengy et al., 2022) and recognized open-source labels (Rehman et al., 2024; Jia et al., 2024).
@@ 5.1.3. Implementation
We implement Slot using Pytorch, comprising approximately 3,000 lines of code. All models operate on Python 3.9, utilizing the Gensim library for the Word2Vec model. Regarding hyperparameters, we set all embedding dimensions to 64, and the learning rate to 0.01. We use Adam as the optimizer, with the similarity loss weight () set at and the regularization weight () set at 0.01. The RL action step size () is configured to 0.02. In Section 11, we provide a detailed discussion on the impact of hyperparameter settings, offering a thorough analysis and explanation of the rationale behind the settings and the experimental logic. All experiments are performed on a server running Ubuntu 18.04 LTS with an Intel 13th Gen Core i7-1360P CPU (12 cores, 16 threads, base frequency 2.20 GHz), 32 GB RAM.
@@ 5.3. Overall Detection Efficacy Comparsion (RQ2)
Bold denotes the best results, and denotes the second-best results.
@@
Obs.2: Comparison in Entity-Level Logs. In entity-level log detection, we compared Slot with other node-level methods. Log2vec treats logs as nodes rather than entities, fragmenting entity information and weakening aggregation. GNN-based methods like MAGIC, Threatrace, and FLASH also failed to address benign behaviors camouflaging malicious entities, making them vulnerable to evasion. In contrast, Slot used reinforcement learning to filter out benign-camouflaged neighbors, resulting in superior detection efficacy compared to the SOTA node-level system, FLASH.
@@ 5.4. System Overhead (RQ3)
Figure LABEL:fig:_Detection_effectiveness shows the impact of batch size on detection effectiveness. Since larger graph batches facilitate learning a more complete local structure in GNN processing, the detection efficiency improves as the batch size increases. However, beyond a certain threshold (The maximum number of entities required to complete most of the system activities), the detection capability stabilizes. Considering that batch size has a significant impact on detection accuracy, we analyzed the CPU and memory overhead under different batch sizes. As shown in Figure LABEL:fig:_RAM_and_CPU, larger batch sizes provide more complex graph structures and node information, which requires more memory and node computation during the graph information propagation process. In terms of detection time, as shown in Figure LABEL:fig:_time_with_batch_size, a larger batch size corresponds to increased detection time. We observe that when the batch size exceeds 5000, the performance improvement in detection becomes marginal, as 5000 nodes are sufficient to ensure that most system behaviors are not fragmented. Under the condition of a batch size of 5000, Figure LABEL:fig:_time illustrates that Slot achieves faster detection speed compared to MAGIC (Jia et al., 2024). Although MAGIC enhances the effectiveness of GNN-based anomaly detection using masking techniques, its KNN distance calculations for anomaly detection still require a significant amount of time.
@@ 5.8. Alert Validation (RQ7)
Case Study.
Using the attack scenario from Section 3.1, we demonstrate how Slot detects APT attacks from audit logs and reconstructs the attack chain. In this scenario, the attacker leverages a Firefox backdoor and embeds benign DNS resolutions to mask their activities. Slot employs semantic and topological embeddings to derive primary node representations and calculate similarities, distinguishing between benign behaviors (e.g., Firefox accessing common IPs) and malicious activities. By segmenting similarities between Firefox1 and Firefox2 and using reinforcement learning, Slot filters out benign activities disguised within attack events. Using GNN, it learns the genuine behavioral patterns of nodes. In the final Attack-chain Reconstruction phase, TTP encoding further separates benign peripheral nodes, resulting in the red-highlighted attack chain. Core malicious activities include communication with IP , downloading and executing a malicious file (), and interacting with IP to manipulate processes for attack scanning.
@@ 6. Conclusion
The persistent and evolving threat posed by advanced persistent threats (APTs) necessitates innovative detection solutions that can navigate the complexities of modern cyberattacks. Existing methods often fall short in addressing challenges like limited detection efficacy, vulnerability to adversarial attacks, and insufficient support for developing defense strategies. To bridge these gaps, we introduced Slot, an APT detection approach leveraging provenance graphs and graph reinforcement learning. Using advanced provenance graph digging, Slot uncovers multi-level hidden relationships among system behaviors. Additionally, it leverages graph reinforcement learning to adapt to new user activities and attack strategies, improving accuracy and resilience in adversarial environments. By automatically constructing attack chains, Slot identifies attack paths with precision and supports defense strategy development. Our evaluations show Slot’s effectiveness, achieving 99% APT detection accuracy and outperforming state-of-the-art approaches. Further experiments highlight Slot’s high efficiency, resilience to adversarial attacks, and ability to support defense strategy development.
@@ Appendix A Appendix
A.2. Comparison of Neighbor Selection in Slot and GAT for Adversarial Attacks
Method | 0 benign event | 5 benign event | 15 benign event | 20 benign event |
Slot | 99% | 98% | 98% | 99% |
GAT | 90% | 85% | 73% | 80% |
The Graph Attention Network (GAT) (Velickovic et al., 2017) utilizes the attention mechanism to compute the weights between nodes, dynamically adjusting the aggregation weights based on the features of neighboring nodes. The core idea of GAT lies in the weighted aggregation of all neighboring nodes, emphasizing the important neighbors while reducing the influence of irrelevant information. Although GAT can handle nodes with different features, it may suffer from misclassification when confronted with masquerading behaviors, particularly when the features of masqueraders and legitimate users are similar. Slot specifically addresses masqueraders by introducing a filtering mechanism that effectively identifies and filters out masquerading neighbors, thereby reducing the impact of masqueraders on the aggregation results. Through reinforcement learning, Slot can continuously optimize the neighbor selection process, enhancing its resistance to masquerading behaviors.
A.3. Impact of Oversampling Ratio

### v4

@@ Abstract.
Advanced Persistent Threats (APTs) represent sophisticated cyberattacks characterized by their ability to remain undetected within the victim system for extended periods, aiming to exfiltrate sensitive data or disrupt operations. Existing detection approaches often struggle to effectively identify these complex threats, construct the attack chain for defense facilitation, or resist adversarial attacks. To overcome these challenges, we propose Slot, an advanced APT detection approach based on provenance graphs and graph reinforcement learning. Slot excels in uncovering multi-level hidden relationships, such as causal, contextual, and indirect connections, among system behaviors through provenance graph mining. Slot implements semi-supervised learning with limited labels through efficient label similarity computation, significantly enhancing both detection performance and model robustness. By pioneering the integration of graph reinforcement learning, Slot dynamically adapts to new user activities and evolving attack strategies, enhancing its resilience against adversarial attacks. Additionally, Slot automatically constructs the attack chain according to detected attacks with clustering algorithms, providing precise identification of attack paths and facilitating the development of defense strategies. Evaluations with real-world datasets demonstrate Slot’s outstanding accuracy, efficiency, adaptability, and robustness in APT detection, with most metrics surpassing state-of-the-art methods. Additionally, case studies conducted to assess Slot’s effectiveness in supporting APT defense further establish it as a practical and reliable tool for cybersecurity protection.
@@ 1. Introduction
To effectively combat APTs, it is crucial to detect them early with both timeliness and precision. Traditional detection methods may struggle with these complex and prolonged attacks, often resulting in either inadequate accuracy or excessive time to produce detection results (Feng et al., 2025, 2023). Recently, researchers have increasingly utilized provenance graphs to trace APT activities, progressively achieving practical efficacy and efficiency.
@@
To address the aforementioned gaps, we introduce Slot, an advanced APT detection approach based on provenance graphs and graph reinforcement learning. Unlike existing methods, Slot achieves enhanced APT detection accuracy by leveraging Latent Behavior Mining and Graph Reinforcement Learning to deeply excavate multi-level hidden relationships (e.g., causal dependencies, contextual interactions, and indirect attack traces) from system behavior data. Additionally, by incorporating graph reinforcement learning, Slot can autonomously learn and adapt to new user activities and attack strategies without relying on threat reports derived from expert insights, thereby enhancing its adaptability and resilience against adversarial attacks. Furthermore, Slot can automatically construct the attack chain, accurately identifying the attack path, which simplifies the defense process.
@@
We propose Slot, an accurate method for Advanced Persistent Threat (APT) detection that maintains robustness in adversarial environments. It effectively identifies malicious activities within large-scale audit logs while reconstructing attack chains to support the development of targeted defense strategies.
@@
Slot implements semi-supervised learning with limited labels through efficient label similarity computation, significantly enhancing both detection performance and model robustness.
@@ 2.2. Graph Reinforcement Learning
Reinforcement Learning (RL) has recently achieved success in addressing challenges across various fields (Feng et al., 2020), including robotics (Levine et al., 2018), gaming (Chen et al., 2022), and natural language processing (NLP) (Uc-Cetina et al., 2023). Researchers have found that RL methods enable effective exploration of the topological structures and attribute information of graphs by analyzing key components such as nodes, links, and subgraphs. SUGAR (Sun et al., 2021) employs Q-learning to adaptively select significant subgraphs to represent discriminative information of the graphs. Bachu et al. (Bacciu and Numeroso, 2022) develop perturbation strategies for local explanations of graph data by optimizing multi-objective scores. Other related works (Yuan et al., 2021; Jiang et al., 2018) have demonstrated impressive performances in graph representation learning through RL, proving its efficacy in this domain. But currently, graph reinforcement learning has not yet been attempted in the field of log provenance detection.
@@ 4.2. Graph Construction
Slot constructs a comprehensive system provenance graph using audit data collected from logging infrastructures such as Windows ETW (Corporation, [n. d.]), Linux Audit (lin, [n. d.]), and CamFlow (Pasquier et al., 2018). It adheres to the existing definitions of provenance graphs and employs recent graph preprocessing techniques (Wang et al., 2022; Rehman et al., 2024), abstracting events into subjects, objects, and relationships, as illustrated in Table 2. Here, source nodes like Process/Thread represent subjects, and target nodes like Files or sockets represent objects, with system call events serving as the relationships to form graph . In addition to focusing on system behaviors, Slot further enriches the relationships between behaviors and explores multi-hop latent relations in Section 4.3. Regarding node features, Slot fully considers meaningful semantic information such as process names and command-line parameters for process nodes, file paths for file nodes, network IP addresses and ports for netflow nodes. Slot effectively mines deep semantic patterns in provenance logs by constructing node ”summary sentences” through aggregating semantic attributes (process names, file paths, IPs) and causal events (system calls) from 1-hop neighbors, incorporating timestamp-sorted event sequences with positional encoding to preserve temporal relationships, and leveraging the Word2Vec model (Mikolov, 2013) to learn vector representations of semantic features.
@@ 4.3. Latent Behavior Mining
A complete user behavior or attack campaign typically involves multiple system calls, which manifest as multi-hop node relationships in provenance graphs. To accelerate node information propagation and uncover implicit dependencies between distant nodes, Slot introduces a Latent Relation Mining Module. This module automatically constructs hidden relationships across multi-hop nodes through a relational attention mechanism, prioritizing causal chains and attack-relevant contextual patterns.
@@ 4.4.2. Adaptive Bandit Neighbor Selector
Reward The optimal at layer under relation aims to find the most similar neighbors to the central node. A decrease in the average neighbor distance signifies that the selected neighbors exhibit higher similarity to the central node. We design a binary reward based on the difference in average distances between two consecutive epochs. We define the reward for epoch as:
@@ 5.1.1. Datasets
The evaluation of Slot was carried out using three open-source datasets: StreamSpot (The, [n. d.]), Unicorn Wget (Han et al., 2020) and DARPA E3 (DAR, [n. d.]). Both StreamSpot and Unicorn consist of batch audit logs generated in controlled environments. The StreamSpot dataset includes 600 batch logs monitoring system calls across six distinct scenarios, five of which simulate benign user behavior, while one represents a drive-by download attack. The Unicorn dataset contains 150 batch logs collected via Camflow (Pasquier et al., 2018), with 125 logs corresponding to benign activity and 25 involving supply chain attacks. These attacks are stealthy in nature, designed to mimic normal system workflows, and are challenging to detect. The DARPA dataset is part of the DARPA Transparent Computing program, gathered during adversarial engagements in enterprise networks. The Red Team conducted APT attacks using various vulnerabilities to exfiltrate sensitive information, while the Blue Team focused on identifying these attacks through network host audits and causal analysis. We conducted experimental comparisons using datasets commonly shared across multiple SOTA algorithms (Wang et al., 2022; Rehman et al., 2024; Jia et al., 2024; Zengy et al., 2022) and recognized open-source labels (Rehman et al., 2024; Jia et al., 2024).
@@ 5.1.3. Implementation
We implement Slot using Pytorch, comprising approximately 3,000 lines of code. All models operate on Python 3.9, utilizing the Gensim library for the Word2Vec model. Regarding hyperparameters, we set all embedding dimensions to 64, and the learning rate to 0.01. We use Adam as the optimizer, with the similarity loss weight () set at and the regularization weight () set at 0.01. The RL action step size () is configured to 0.02. In Section 5.6, we provide a detailed discussion on the impact of hyperparameter settings, offering a thorough analysis and explanation of the rationale behind the settings and the experimental logic. All experiments are performed on a server running Ubuntu 18.04 LTS with an Intel 13th Gen Core i7-1360P CPU (12 cores, 16 threads, base frequency 2.20 GHz), 32 GB RAM.
@@ 5.3. Overall Detection Efficacy Comparsion (RQ2)
Bold denotes the best results, and denotes the second-best results.
represents the percentage of decrease, represents the percentage of improvement.
@@
Improvement compared to the average of all detection systems except Slot.
2
FPR means False Positive Rate. A detection system with a lower false positive rate can better avoid false positive attacks during detection.
@@
Obs.2: Comparison in Entity-Level Logs. In entity-level log detection, we compared Slot with other node-level methods. Log2vec treats logs as nodes rather than entities, fragmenting entity information and weakening aggregation. GNN-based methods like MAGIC, Threatrace, and FLASH also failed to address benign behaviors camouflaging malicious entities, making them vulnerable to evasion. In contrast, Slot used reinforcement learning to filter out benign-camouflaged neighbors, resulting in superior detection efficacy compared to the SOTA node-level system, FLASH. In Figure 6, compared to FLASH, the Slot ROC curve demonstrates superior model classification performance.
@@ 5.4. System Overhead (RQ3)
| Phase | Graph Construction | Reinforcement Learning | GNN |
| Time cost | 46.12s | 331.94s | 64.69s |
| Percentage | 10.41% | 74.97% | 14.67% |
@@
Figure LABEL:fig:_Detection_effectiveness shows the impact of batch size on detection effectiveness. Since larger graph batches facilitate learning a more complete local structure in GNN processing, the detection efficiency improves as the batch size increases. However, beyond a certain threshold (The maximum number of entities required to complete most of the system activities), the detection capability stabilizes. Considering that batch size has a significant impact on detection accuracy, we analyzed the CPU and memory overhead under different batch sizes. As shown in Figure LABEL:fig:_RAM_and_CPU, larger batch sizes provide more complex graph structures and node information, which requires more memory and node computation during the graph information propagation process. In terms of detection time, while the introduction of reinforcement learning incurs additional overhead, as shown in Table 7, our tests on graphs with 300K nodes demonstrate that although reinforcement learning consumes more time, the overhead can be considered acceptable.
@@ 5.8. Alert Validation (RQ7)
Case Study. Using the attack scenario from Section 3.1, we demonstrate how Slot detects APT attacks from audit logs and reconstructs the attack chain. In this scenario, the attacker leverages a Firefox backdoor and embeds benign DNS resolutions to mask their activities. Slot employs semantic and topological embeddings to derive primary node representations and calculate similarities, distinguishing between benign behaviors (e.g., Firefox accessing common IPs) and malicious activities. By segmenting similarities between Firefox1 and Firefox2 and using reinforcement learning, Slot filters out benign activities disguised within attack events. Using GNN, it learns the genuine behavioral patterns of nodes. In the final Attack-chain Reconstruction phase, TTP encoding further separates benign peripheral nodes, resulting in the red-highlighted attack chain. Core malicious activities include communication with IP , downloading and executing a malicious file (), and interacting with IP to manipulate processes for attack scanning.
5.9. Discussion
[…]
@@ 6. Conclusion
The persistent and evolving threat posed by advanced persistent threats (APTs) necessitates innovative detection solutions that can navigate the complexities of modern cyberattacks. Existing methods often fall short in addressing challenges like limited detection efficacy, vulnerability to adversarial attacks, and insufficient support for developing defense strategies. To bridge these gaps, we introduced Slot, an APT detection approach leveraging provenance graphs and graph reinforcement learning. Using advanced provenance graph digging, Slot uncovers multi-level hidden relationships among system behaviors. Slot boosts semi-supervised learning with limited labels via efficient label similarity computation, enhancing both detection accuracy and model robustness. Additionally, it leverages graph reinforcement learning to adapt to new user activities and attack strategies, improving accuracy and resilience in adversarial environments. By automatically constructing attack chains, Slot identifies attack paths with precision and supports defense strategy development. Our evaluations show Slot’s effectiveness, achieving 99% APT detection accuracy and outperforming state-of-the-art approaches. Further experiments highlight Slot’s high efficiency, resilience to adversarial attacks, and ability to support defense strategy development.
7. Acknowledgments
This research is funded by the National Key Research and Development Program of China (2023YFE0111100), Natural Science Basic Research Program of Shaanxi (No. 2025JC-JCQN-073), National Natural Science Foundation of China under Grant (No. 62272370), Young Elite Scientists Sponsorship Program by CAST (2022QNRC001), the China 111Project (No.B16037), Qinchuangyuan Scientist + Engineer Team Program of Shaanxi (No. 2024QCY-KXJ-149), Songshan Laboratory (No. 241110210200), Open Foundation of Key Laboratory of Cyberspace Security, Ministry of Education of China (No.KLCS20240405 ) and the Fundamental Research Funds for the Central Universities (QTZX23071), the National Research Foundation, Singapore, and DSO National Laboratories under the AI Singapore Programme (AISG Award No: AISG2-GC-2023-008), the National Research Foundation, Singapore, and the Cyber Security Agency under its National Cybersecurity R&D Programme (NCRP25-P04-TAICeN), the National Research Foundation, Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) programme, and Ripple under its University Blockchain Research Initiative (UBRI) (Feng et al., 2022).
@@ Appendix A Appendix
A.2. Impact of Oversampling Ratio